40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 795     Entry time: Wed Jun 2 20:24:26 2010
Author: rana, alastair 
Type: Misc 
Category: GYRO 
Subject: Optics moved 

Today we moved some of the optics away from the blue box to make room for the mode matching.

We also found a 1" diameter optic in one of the cabinets with a transmission of 0.81 % (S-pol). This makes a nice match with the 0.57% transmission of the 2" diameter output coupler and so its now installed. We'll later replace it with a good 2" diameter mirror.

We also moved all of the optics after the laser to align with the screw holes and are trying to make the beams go along the screw hole lines.

Alastair has remeasured the laser beam profile after the EOM and the PBSs using the WinCamD beam scanner. Results are being produced.

ELOG V3.1.3-